OpenStudy (anonymous):

Simplify: (2x + 8)(2x - 8). A. 4x2 - 16 B. 4x2 - 64 C. 4x2 - 16x - 64 D. 4x2 - 8x + 16 I think its C.

OpenStudy (anonymous):

Or A

hero (hero):

Hint: (a+b)(a-b) = a^2 - b^2

OpenStudy (anonymous):

So C . correct ?

hero (hero):

2nd Hint: (2x+8)(2x-8) = (2x)^2 - (8)^2

OpenStudy (anonymous):

D

hero (hero):

3rd Hint: (2x)^2 - (8)^2 = (2*2*x*x) - (8*8)